OpenStudy (anonymous):

From the data below, calculate the total heat (in J) needed to convert 0.602 moles of gaseous ethanol at 300.0 Celsius and 1 atm to liquid ethanol at 25.0 celsius and 1 atm: b.p. at 1 atm 78.5 celsius, ΔHvap: 40.5 Kj/mol, Cethanol (g) 1.43 j/g c, Cethanol(l) 2.45 j/g c

OpenStudy (aaronq):

 \(q_3=m*C_{(g)}*\Delta T\) \(q_2=\Delta H_{vap}*n_{Et}\) \(q_1=m*C_{(l)}*\Delta T\) \(T_q=q_1+q_2+q_3\) Note that the heat capacities given are in "\(J/g*C^{\;o}\)" and \(\Delta H_{vap}\) in "kJ/mol".